OpenStudy (anonymous):

You and your friends decide to have a submarine sandwich sale as a fundraiser for the school band. You spend $83 on supplies and charge $5 per sandwich. At the end of the day your profit is $962. How many sandwiches did you and your friend sell? Write an equation.

hero (hero):

83+ 5x = 962 962 - 83 = 5x 879 = 5x 879/5 = x 175.8 or about 176 sandwiches